You are dealt one card from a 52-Card deck. Calculate the probability of and odds for being dealt an eight or a diamond.

The probability is ____

The odds for being dealt and eight or diamond are _____.
Answer #1

P(8 or diamond) = P(8) + P(diamond) - P(8 and diamond) = 4/52 + 13/52 - 1/52 = 16/52 = 4/13

The probability is 4/13

The odds are = 16 / (52 - 16) = 16/36 = 4/9